Two kinds of bacteria are found in a sample of tainted food. It is found that the populations size of type 1,\\(N_1\\) and of type 2, \\(N_2\\) satisfy the equations \\(\\frac{dN_1}{dt}=-k_1N_1, N_1(0)=N_{1,0}\\) and \\(\\frac{dN_2}{dt}=-k_2N_2, N_2(0)=N_{2,0}\\) . Then the population sizes equal \\(N_1=N_2\\) at the following time
1
Expert's answer
2020-03-19T16:14:42-0400

"\frac{dN_1}{dt}=-k_1N_1, \quad N_1(0)= N_{1,0}"


"\frac{dN_2}{dt}=-k_2N_2, \quad N_2(0)= N_{2,0}"

We have the following solutions:

"N_1(t)=N_{1,0}e^{-k_1t}" and "N_2(t)=N_{2,0}e^{-k_2t}"

Suppose that at the time "t=t_0" the population sizes are equal:

"N_1(t_0)=N_2(t_0)"

"N_{1,0}e^{-k_1t_0} =N_{2,0}e^{-k_2t_0}"

"\frac{N_{1,0}}{N_{2,0}}=e^{(k_1-k_2)t_0}"

"(k_1-k_2)t_0=\ln(\frac{N_{1,0}}{N_{2,0}})"

"t_0=\frac{1}{k_1-k_2} \ln(\frac{N_{1,0}}{N_{2,0}})"

Answer: at the time "t_0=\frac{1}{k_1-k_2} \ln(\frac{N_{1,0}}{N_{2,0}})"